OpenStudy (anonymous):

PLEASE: Simplify...(9X)^-2

OpenStudy (anonymous):

1/9x^2

OpenStudy (anonymous):

the negative makes it a fraction

OpenStudy (anonymous):

By the Law of Exponents: \[(ab)^n = (a)^n (b)^n\] Thus, \[(9x)^-2 = (9)^-2 (x)^-2\] 9^-2 = 1/81 x^-2 = 1/x2

OpenStudy (anonymous):

Ok, so I write it as a fraction? the answer is:1/9x^2 ??

OpenStudy (anonymous):

yupsiieeesss

OpenStudy (anonymous):

or 1/81x^2

OpenStudy (anonymous):

1/81x^2...this is a fraction where 1 is placed over the 81 with 81 having an exponent of 2?

OpenStudy (anonymous):

I am confused...is the correct answer 1/9x^2 or is the correct answer 1/81x^2

OpenStudy (anonymous):

both LOL